Craig is 18 years old and his father is 46. Solve an equation to find in how many years’ time Craig’s father will be three times as old as Craig.
Answer by Edwin McCravy(20056)

Craig is 18 years old and his father is 46. Solve an equation to find in how many years’ time Craig’s father will be three times as old as Craig.
That's impossible, because the older any two people get, the closer the ratio of
their ages gets to 1.  Their present ratio 46/18 is about 2.6, so the ratio
will only get closer to 1, and will never be 3. 

So the question should be corrected to read:

how many years AGO was Craig’s father three times as old as Craig.
Craig's age y years ago = 18-y
Father's age y years ago = 46-y

46-y = 3(18-y)

46-y = 54-3y
  2y = 8
   y = 4

So 4 years ago, Craig was 14 and his father was 46-4 = 42.

Edwin